OpenStudy (anonymous):

solve ......2/3-6h=-13/6

OpenStudy (anonymous):

Is this the question ? \[\frac{2}{3}-6h=-\frac{13}{6}\]

OpenStudy (anonymous):

subtract the 2/3 from both sides and then devide by -6 from both sides

OpenStudy (anonymous):

h= -17/36..got tis by solving the above eqn..cross check by substituting the value of h in the above eqn..u l get -13/6..